OpenStudy (anonymous):

Find the equation of the plane tangent to f(x,y)= x^2+2xy^2+x^3y when x=2 and y=3 Calculus 3

OpenStudy (anonymous):

Start by partially differentiating your function with respect to x, and then with respect to y.

OpenStudy (anonymous):

fx=2x+2y^2+x^3y fy=4xy+x^3 then what?

OpenStudy (anonymous):

Ok so now you put your values of (2,3) into each one of those

OpenStudy (anonymous):

fx=58 and fy=32

OpenStudy (anonymous):

i got an answer of z-64=58(x-2)+32(y-3) but that does not sound correct.

OpenStudy (anonymous):

Ooh, you're one step ahead of me, I'm doing at the same time. Yeah that's what I get, and I think it's right, well done :)

OpenStudy (anonymous):

Just multiply those brackets out to get z = 58x +32y -148

OpenStudy (anonymous):

Ok! Thank you very much!

OpenStudy (anonymous):

No problem :)